OpenStudy (anonymous):

solve 2x - 4/ x +1 > 0

OpenStudy (mathmale):

You may simplify this problem for yourself and others if you'd please use ( ) to group terms in the numerator and the den. together. For example, do you mean 2x - 4/x +1 > 0, or do you mean (2x-4) / (x+1) ? This is really important as you try to convey math symbolism accurately to others.

OpenStudy (anonymous):

2x - 4/x +1 > 0,

OpenStudy (anonymous):

the first equation u mentioned

OpenStudy (phi):

it it this \[ \frac{2x-4}{x+1} > 0 \]?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

exactly as username "phi" has it

OpenStudy (mathmale):

So, those parentheses make all the difference in the world. Looking at phi's inequality: for which x value is the rational function undefined? For which x value is the numerator of this fraction = to 0?

OpenStudy (phi):

ok, that means if you type it in, you should use parens, and it would look like this (2x-4) / (x+1) > 0 Or , use the equation editor (see the button on the lower left of where you input )

OpenStudy (anonymous):

\[2x-\frac{4}{x}+1>0\] or \[\frac{2x-4}{x+1}>0\]

OpenStudy (anonymous):

the second equation tomhue

OpenStudy (phi):

to answer the question, notice that if the top is positive and (at the same time) the bottom is positive, the quotient is positive. And positive means > 0 OR if both the top and bottom are negative (at the same time) the quotient will be positive (meaning > 0)

OpenStudy (phi):

Can you figure out what values make x+1 > 0 true ?

OpenStudy (anonymous):

I solved the equation and ended up with x = -1 and x = 2, so now I am trying to write it out as an equation. I know has has to be greater than 2 but it can also be less than -1, like -4...

OpenStudy (phi):

finding x= -1 and x=2 is helpful. But we have to think about it. First, do you agree that if the top number is positive and the bottom number is positive , then the result will be positive ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

and do you agree that positive number is > 0 ? I assume you think so. you found that the top 2x-4=0 when x is 2 if x is bigger than 2, then 2x-4 will be bigger than 0. i.e. 2x-4 will be positive. the bottom x+1 will also be positive in other words, if x>2 then (2x-4) is positive and x+1 is positive and the result is positive . one way for (2x-4)/(x+1) > 0 to be true is for x > 2

OpenStudy (anonymous):

ok makes sense

OpenStudy (phi):

you also found x+1=0 when x = -1 if x is less than -1, x+1 will be negative. the top 2x - 4 with x < -1 will also be < 0 when the top is negative, and the bottom is negative (2x-4)/(x+1) will be positive (negative divided by negative is positive) so another way for (2x-4)/(x+1) > 0 is when x < -1

OpenStudy (phi):

and just to finish off the thought, if x is bigger than -1 but less than 2 the top will be negative and the bottom will be positive, and the result of negative divided by positive will be negative... meaning (2x-4)/(x+1) < 0 (which we do not want)

OpenStudy (anonymous):

ok, so what is the final answer?

OpenStudy (mathmale):

Hope you won't object to my presenting a different approach, one based on phi's work. Earlier I asked you to tell me what x value makes the numerator = 0, and I also asked you which x value makes the den. = 0. Those values are 2 and -1 respectively. (BRB)

OpenStudy (mathmale):

Graph these critical numbers on a number line as follows. Then identify the intervals defined by those 2 critical numbers: